OpenStudy (pgreene008):

A building has a ramp to its front doors to accommodate the handicapped. If the distance from the building to the end of the ramp is 19 feet and the height from the ground to the front doors is 4 feet, how long is the ramp? (Round to the nearest tenth.)

OpenStudy (pgreene008):

@ParthKohli @dan815

OpenStudy (pgreene008):

@ brill

OpenStudy (anonymous):

You can set up a diagram like this, I always find that helpful

OpenStudy (anonymous):



OpenStudy (anonymous):

Do you know how to set up Pythagorean theorem?

OpenStudy (pgreene008):

yeah

OpenStudy (anonymous):

It's \[a^2+b^2=c^2\]In this case a=19 and b=4

OpenStudy (anonymous):

Can you solve that for c please?

OpenStudy (pgreene008):

377

OpenStudy (pgreene008):

@wolf1728 can you help me?

OpenStudy (wolf1728):

Sure what is 19 * 19 ?

OpenStudy (pgreene008):

361

OpenStudy (anonymous):

You're close @pgreene008 I think you just did 19^2+4^2 Remember it is c^2 on the other side so it is\[377=c^2\]

OpenStudy (anonymous):

Do you know how to solve for c now?

OpenStudy (pgreene008):

its sqrt(377)

OpenStudy (wolf1728):

Sorry for making you repeat your work. Anyway if c^2 = 377 then c equals ?

OpenStudy (pgreene008):

sqrt(377)

OpenStudy (anonymous):

Yup

OpenStudy (wolf1728):

correct - and the sq root of 377 euqals?

OpenStudy (anonymous):

If you wanted an exact answer then I suggest typing into a calculator Remember your units, feet

OpenStudy (pgreene008):

19.4 feet?

OpenStudy (wolf1728):

That is it

OpenStudy (anonymous):

Yup Although I would double check with your teacher if she wants a decimal approx or an exact answer

OpenStudy (anonymous):

Actually wolf1728 both versions are the same thing In fact sqrt377 might be more correct because it is the exact answer

OpenStudy (wolf1728):

Yes, brill that is true

OpenStudy (pgreene008):

Thanks for the help guys!!! :) @wolf1728 @Brill

OpenStudy (wolf1728):

u r welcome

OpenStudy (anonymous):

No problem :)